OpenStudy (isabel☺):

how to solve (d^(2) y/dx^2) + k^(2) y = 0??

OpenStudy (anonymous):

let y=Acosx+Bsinx

OpenStudy (isabel☺):

DE?

OpenStudy (anonymous):

sorry I responded but somehow posting is missing

OpenStudy (anonymous):

you can start by writing down the characteristics equation

OpenStudy (anonymous):

m^2+k^2=0 so m=+-ki therefore y=Acoskx+Bsinkx I left out the k in the initial response

OpenStudy (isabel☺):

okay..

OpenStudy (isabel☺):

i just don't know how :(

ganeshie8 (ganeshie8):

which part is confusing ?

OpenStudy (isabel☺):

i don't know how to solve DE like those.

ganeshie8 (ganeshie8):

So basically you want to solve a second order linear homogeneous DE of form : \(\large y'' + Ay' + By = 0\)

OpenStudy (isabel☺):

yes.. please.. :)

ganeshie8 (ganeshie8):

Asume the required solution looks like below : \(\large y = c_1 y_1 + c_2 y_2\)

ganeshie8 (ganeshie8):

where \(c_1\) and \(c_2\) are arbitrary constants, \(y_1\) and \(y_2\) are two different solutions to the DE.

ganeshie8 (ganeshie8):

So our goal is to simply find two solutions \(y_1\) and \(y_2\) then plug them in above ^^

ganeshie8 (ganeshie8):

Okay, so far ? :)

ganeshie8 (ganeshie8):

Notice that two arbitrary constants occur because it is a second order DE.

OpenStudy (isabel☺):

may you show me your solution please>?

ganeshie8 (ganeshie8):

solution is easy : \(\large y = c_1 \sin (kt) + c_2 \cos (kt) \)

OpenStudy (isabel☺):

and then?

ganeshie8 (ganeshie8):

thats the solution ^^

OpenStudy (isabel☺):

ohh!! hihi thanks! wait what's the difference bet. ORDER and DEGREE of a DE?

ganeshie8 (ganeshie8):

order = highest `derivative` in the DE

ganeshie8 (ganeshie8):

degree = power of the highest derivative

ganeshie8 (ganeshie8):

for example : \[\large \left(\dfrac{d^3y}{dx^3}\right)^2 + \dfrac{dy}{dx} = 1\]

ganeshie8 (ganeshie8):

order = 3 degree = 2

OpenStudy (isabel☺):

ahhhhhh!!!! THANKS!! ☺